# Multi-Domain Evolutionary Optimization of Network Structures

### v1

@@ Abstract
Multi-Task Evolutionary Optimization (MTEO), an important field focusing on addressing complex problems through optimizing multiple tasks simultaneously, has attracted much attention. While MTEO has been primarily focusing on task similarity, there remains a hugely untapped potential in harnessing the shared characteristics between different domains to enhance evolutionary optimization. For example, real-world complex systems usually share the same characteristics, such as the power-law rule, small-world property and community structure, thus making it possible to transfer solutions optimized in one system to another to facilitate the optimization. Drawing inspiration from this observation of shared characteristics within complex systems, we set out to extend MTEO to a novel framework—multi-domain evolutionary optimization (MDEO). To examine the performance of the proposed MDEO, we utilize a challenging combinatorial problem of great security concern—community deception in complex networks as the optimization task. To achieve MDEO, we propose a community-based measurement of graph similarity to manage the knowledge transfer among domains. Furthermore, we develop a graph representation-based network alignment model that serves as the conduit for effectively transferring solutions between different domains. Moreover, we devise a self-adaptive mechanism to determine the number of transferred solutions from different domains and introduce a novel mutation operator based on the learned mapping to facilitate the utilization of knowledge from other domains. Experiments on eight real-world networks of different domains demonstrate MDEO’s superiority in efficacy compared to classical evolutionary optimization. Simulations of attacks on the community validate the effectiveness of the proposed MDEO in safeguarding community security.
@@
Complex network, multi-domain evolutionary optimization, multi-task evolutionary optimization, transfer learning.
@@ I Introduction
Multi-task evolutionary optimization (MTEO) has stood as an emerging framework to solve complex optimization problems, undergoing remarkable advancements [1, 2, 3, 4] since the multi-factorial evolutionary algorithm (MFEA) was proposed [5]. MTEO aims to address situations where there are multiple related tasks or objectives being optimized simultaneously, with the goal of achieving better performance on each task.
@@
While most existing endeavors primarily concentrate on exploring relationships between tasks, a research gap persists regarding correlations between different domains in evolutionary optimization. In real-world applications, we encounter a multitude of networks that capture different aspects of complex systems [6]. These networks may represent social networks [7] or transportation networks [8]. Many of these networks share common structural properties, such as power-law distribution [9, 10], community structure [11, 12] and small-world characteristic [13, 14, 15].
@@
In real-world scenarios, different domains often have common objective, such as the task of critical node mining in complex networks [16, 17], which can contribute to the prevention of catastrophic outages in the power networks [18], the identification of drug target candidates in protein networks [19], or the improvement of the robustness in communication networks [20]. The existence of shared attributes and structural properties across real-world networks presents an avenue for knowledge transfer and collaborative learning when optimizing an identical task. To date, multi-task evolutionary optimization of complex networks has already been extensively studied [21, 22], however, leveraging the correlation of domains to facilitate the optimization of the network structure remains underexplored. Inspired by MTEO, we propose a new framework called multi-domain evolutionary optimization (MDEO) for network structures, in which solutions obtained from optimizing one network can be effectively transferred and adapted to optimize another network with improved efficacy. The following are the main differences between MTEO and MDEO, justifying the novelty of our work:
@@
Granularity of Tasks: In MTEO, the emphasis is on optimizing different tasks, which may be distinct and well-defined within a single network. Conversely, MDEO focuses on optimizing an identical task in multiple networks that may represent different complex systems. The goal of MDEO is to find the optimal solution for each domain.
Knowledge Transfer Across Networks: The solutions aiming at different tasks are aligned since MTEO is within the same domain. In the case of MDEO, the nodes or edges in one network are different from their counterparts in another network with the same index.
@@
In this study, we take a further stride by delving into the case of more than three networks, i.e., many networks will be involved in the optimization. The proposed MDEO consists of four components 1) graph similarity 2) graph embedding 3) network alignment 4) many-network evolutionary optimization and its overview is as follows:
@@
Multi-domain evolutionary optimization would present a significant challenge due to the inherent computational complexity when dealing with a large number of networks. Additionally, the efficacy of solutions on one network may not be necessarily translated to success on others. To address these issues, we propose a new measurement of graph similarity, quantifying the closeness between networks at the community level. The calculation of similarity enables us to focus on a subset of closely related networks during the knowledge transfer process, thereby effectively reducing the computational burden associated with multi-domain optimization. Moreover, we are able to prioritize the solutions from those networks with a higher degree of similarity based on the proposed graph similarity, thereby increasing the likelihood of successful adaptations.
In the multi-domain context, transferring solutions across different networks necessitates the establishment of node correspondences. We first employ a graph autoencoder to obtain graph embeddings aimed at capturing node similarity and higher-order interactions.
@@
Building upon the derived graph embeddings, we then propose a novel network alignment method that combines supervised and unsupervised learning. In supervised learning, we propose a community-based anchor node selection method to build up the training set and improve the alignment accuracy. This approach empowers us to create mappings of nodes sharing analogous roles or structures across different networks.
@@
In the process of many-network evolutionary optimization, we transfer solutions from highly similar networks to the target network, achieved by the utilization of the node mappings obtained through network alignment. By observing the contributions of these transferred solutions, we propose a self-adaptive model to determine the appropriate number of solutions to transfer between networks, with the consideration of the previously calculated graph similarity. Additionally, we develop a cross-domain mutation operator that makes better use of the knowledge from other networks to enhance the optimization.
In the validation with eight real-world networks, the experimental results reveal that the average fitness achieved within the MDEO framework outperforms that of classical evolutionary optimization, underscoring the effectiveness of our proposed framework. With the simulations of different attacks on the community, the results of metrics of normalized mutual information (NMI) [23] and adjusted rand index (ARI) [24] reveal that our proposed model can safeguard the community information from being exploited improperly.
@@
In this work, we utilize the term ‘graph’ in the computational context and employ ‘network’ to denote real-world systems. Both terms refer to data composed of nodes and edges. The subsequent sections of the paper are organized as follows: Section II provides an overview of relevant literature, and Section III introduces the optimization problem–community deception. Section IV delves into the intricacies of our proposed MDEO framework, detailing its methodology and components. Section V entails an examination of the effectiveness of our method through a series of experiments conducted on various real-world networks. Lastly, Section VI concludes with a summary of our main results.
II Related works
@@
The wide applicability of evolutionary algorithms in handling discrete problems with non-linear characteristics has led to their extensive utilization within complex networks. For example, Wang et al. [25] develop a multi-objective model by optimizing the computational complexity and nonuniform latency to enhance the tolerance of networks against attacks. Similarly, Wu et al. leverage the inherent community structure of networks to downsize the decision space and introduce a multi-objective framework for network reconstruction by considering the reconstruction error and sparsity in [26]. Based on the same optimization objectives, Ying et al. propose a logistic principal component analysis-based method to improve the representation of networks in [27].
@@
Meanwhile, evolutionary algorithms have also been applied to address the problem of community deception. Chen et al. use the genetic algorithm to minimize the modularity after the modification on edges in [28]. Further, they categorize community deception based on the scales as community deception on node communities, target community and the entire community structure in [29]. A self-adaptive evolutionary framework is then proposed to facilitate the search of the optimal edge set and a permanence-based method is developed to reduce the search space in [30]. To alleviate the problem of the evaluation expense, Zhao and Cheong [31] propose a divide-and-conquer strategy to split the graph into subgraphs and optimize them with the genetic algorithm individually. Moreover, evolutionary algorithms have been applied to various other aspects of complex networks, including dynamic community detection [32], and module identification [33]. These applications demonstrate the versatility and effectiveness of evolutionary algorithms in tackling diverse challenges within the field of complex network analysis.
@@
To date, considerable efforts have been directed towards tackling multi-task problems using evolutionary algorithms. In [5], Gupta et al. propose a paradigm, multi-factorial evolutionary algorithm, which can optimize different tasks simultaneously and has laid the foundation for the related research. For example, the optimization on single objective has been extended to the case of multiple objectives in [34], adopting the implicit transfer mode where the crossover is operated on the solutions aiming at different tasks [35, 36]. There is another route referred to explicit transfer that deliberately incorporates task-specific information to guide knowledge exchange between tasks by using statistic information [37] or constructing a mapping matrix [38, 39].
@@
Beyond traditional fields, MTEO has been exploited within complex networks, such as in the optimization of network robustness. In [22], Wang et al. utilize the correlation of tasks to develop a new crossover operator, thereby obtaining knowledge from different individuals. As indicated in [40], structural destruction and cascading failure, two seemingly independent objectives, can promote each other in the optimization by developing a graph neural network-based method to exchange information between solutions aiming at different tasks. In one study of influence maximization [41], each task is treated as a transformation by formula, enabling the application of MTEO in this problem. In [21], Wu et al. promote the reconstruction of multiplex networks using MTEO with the inherent correlation of different layers. Similar work can be found in [42], in which the optimization of modularity in each layer is modeled as an individual task, thereby facilitating the community detection in multiplex networks.
Despite that networks often share the same characteristics, cross-domain evolutionary optimization, especially concerning network structures, has received limited attention. Motivated by this overlooked potential, we introduce a novel optimization framework—Multi-Domain Evolutionary Optimization.
@@ III Optimization problem formulation
Community deception involves nodes intentionally concealing their true community affiliations by modifying (adding and removing) connections. In this study, we select this challenging combinatorial problem [43, 44] in complex networks as an illustration of our MDEO for the following concerns:
Exploiting Community Detection Algorithms: With the assistance of community detection algorithms, densely interconnected groups of nodes with high intra-group connectivity and low inter-group connectivity can be easily identified [42, 45]. The development of community detection, while useful for understanding network structures and interactions, can also have certain dangerous aspects caused by the potential misuse or exploitation of community information, as suggested in [46, 47].
@@
Community deception serves as an innovative advancement in network science, radically altering node affiliations through minimal topological adjustments. Specifically, the modification on the network involves the addition of nonexistent edges and the deletion of existing edges. Let $G=\{V,E\}$ be a network where $V$ and $E$ denote the nodes and edges, respectively. The updated edge set can be defined as follows:
$E^{\prime}=\left(E\cup E^{+}\right)\backslash E^{-},$ (1)
where $E^{+}$ and $E^{-}$ refers to the addition and deletion of edges, that is:
@@
where $\beta$ refers to the perturbance budget, meaning the number of edges that can be modified.
Assume $\widetilde{\mathcal{C}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}^{\prime}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$ are the community partitions of original network $G$ and modified network $G^{\prime}$, then community deception can be represented as a combinatorial problem:
@@
The function $\phi$ refers to the difference between the community structures, before and after modification.
IV Multi-Domain evolutionary optimization on community deception
In this section, the proposed MDEO will be introduced in detail, and its diagram can be found in Figure 1.
Figure 1: The diagram of multi-domain evolutionary optimization. Four networks representing different systems are optimized simultaneously. The transfer of knowledge across networks is achieved through the trained network alignment model. $e^{i}_{j}$ denote the $j$th edge in $i$th network and $M_{i,j}$ is the edge mapping from $i$th network to $j$th network. The solution with the colors of yellow and light red refers to the elite solution and transferred solution from other networks, respectively. The red edge is the mutated edge that will be replaced by the edge of elite solutions of other networks.

### v2

# Multi-Domain Evolutionary Optimization on Combinatorial Problems in Complex Networks

@@ Abstract
Knowledge transfer-based evolutionary optimization has garnered significant attention, such as in multi-task evolutionary optimization (MTEO), which aims to solve complex problems by simultaneously optimizing multiple tasks. While this emerging paradigm has been primarily focusing on task similarity, there remains a hugely untapped potential in harnessing the shared characteristics between different domains. For example, real-world complex systems usually share the same characteristics, such as the power-law rule, small-world property and community structure, thus making it possible to transfer solutions optimized in one system to another to facilitate the optimization. Drawing inspiration from this observation of shared characteristics within complex systems, we present a novel framework, multi-domain evolutionary optimization (MDEO). First, we propose a community-level measurement of graph similarity to manage the knowledge transfer among domains. Furthermore, we develop a graph learning-based network alignment model that serves as the conduit for effectively transferring solutions between different domains. Moreover, we devise a self-adaptive mechanism to determine the number of transferred solutions from different domains, and introduce a knowledge-guided mutation mechanism that adaptively redefines mutation candidates to facilitate the utilization of knowledge from other domains. To evaluate its performance, we use a challenging combinatorial problem known as adversarial link perturbation as the primary illustrative optimization task. Experiments on multiple real-world networks of different domains demonstrate the superiority of the proposed framework in efficacy compared to classical evolutionary optimization.
@@
Complex network, multi-domain evolutionary optimization, knowledge transfer, combinatorial problem.
@@ I Introduction
Knowledge transfer in evolutionary optimization has emerged as a promising paradigm for tackling complex optimization problems by enabling the sharing of useful information across related problems. One representative framework is multi-task evolutionary optimization (MTEO), which has seen notable progress in recent years [1, 2], following the introduction of the multi-factorial evolutionary algorithm (MFEA) [3]. MTEO is designed to handle scenarios involving multiple related tasks being optimized concurrently, with the aim of leveraging inter-task similarity to improve performance on each task [4, 5].
I-A Motivation
@@
In this work, we shift the focus from task-based settings to the structural properties of the network: While most existing endeavors primarily concentrate on exploring relationships between tasks, a research gap persists regarding correlations between different domains in evolutionary optimization. In real-world applications, we encounter a multitude of networks representing different complex systems [6]. These networks may represent social networks [7] or transportation networks [8], many of which share common structural properties, such as power-law distribution [9, 10], community structure [11, 12] and small-world characteristic [13, 14, 15]. In real-world scenarios, the same objective often exists in different domains, such as the task of critical node mining in complex networks [7, 16, 17], which can contribute to the prevention of catastrophic outages in power networks [18], the identification of drug target candidates in protein networks [19], or the improvement of the robustness in communication networks [20]. The existence of shared attributes and structural properties across real-world networks presents an avenue for knowledge transfer and collaborative optimization when dealing with an identical task across different systems.
@@
To date, MTEO in complex networks has already been extensively studied [21, 22], however, leveraging the correlation of domains to facilitate the optimization of network structure remains underexplored. Therefore, we propose a new framework called multi-domain evolutionary optimization (MDEO) for network-structured combinatorial problems, in which solutions obtained from optimizing one network can be effectively transferred and adapted to optimize another network with improved efficacy. The following are the main differences between MTEO and MDEO, justifying the novelty of our work:
@@
Granularity of Tasks: In MTEO, the emphasis is on optimizing different tasks within a single domain. Conversely, MDEO focuses on optimizing an identical task in multiple networks representing different complex systems. The goal of MDEO is to find the optimal solution for each domain. Note that if multiple tasks concurrently exist across different domains, MDEO could be extended to multi-domain multi-task evolutionary optimization.
Knowledge Transfer Across different domains: In MDEO, knowledge is transferred across distinct domains, such as the social network, power network and biology network, enhancing solutions by leveraging insights from one domain to another. This contrasts with MTEO, where knowledge is transferred within the same domain across tasks with similar characteristics, whether continuous or combinatorial.
I-B Contributions
@@
In this study, we take a further stride and delve into the case of more than three networks, i.e., many networks will be involved in the optimization. The proposed MDEO consists of four components 1) graph similarity; 2) graph embedding; 3) network alignment; 4) many-network evolutionary optimization and its overview is as follows:
@@
Multi-domain evolutionary optimization may present a significant challenge due to the inherent computational complexity when dealing with a large number of networks. Additionally, the efficacy of solutions on one network may not necessarily translate to success on others. To address these issues, we propose a new measurement of graph similarity, quantifying the closeness between networks at the community level. The similarity measurement enables us to focus on a subset of closely related networks during the knowledge transfer process, thereby effectively reducing the computational burden and increasing the likelihood of successful adaptations.
@@
In the multi-domain context, transferring solutions across different networks necessitates the establishment of node correspondences. We first employ the graph autoencoder to obtain graph embeddings aimed at capturing node similarity and higher-order interactions. Building upon the derived graph embeddings, we then propose a novel network alignment method that combines supervised and unsupervised learning. In supervised learning, we propose a community-level anchor node selection method to build up the training set and improve the alignment accuracy. This approach empowers us to create mappings of nodes sharing analogous roles or structures across different networks.
@@
In the process of many-network evolutionary optimization, we transfer solutions from similar networks to the target network, achieved by the node mappings obtained through network alignment. By observing the contributions of these transferred solutions, we leverage a self-adaptive model to determine the appropriate number of solutions to transfer between networks, with the consideration of the previously calculated graph similarity.
Validation on eight real-world networks for edge-level tasks reveals that the MDEO framework achieves higher average fitness than classical evolutionary optimization. This trend is also observed in the node-level tasks, indicating the generalizability and adaptability of MDEO.
I-C Organization
@@
The subsequent sections of the paper are organized as follows: Section II provides an overview of related work, and Section III introduces the optimization problem–community deception. Section IV delves into the intricacies of our proposed MDEO framework, detailing its methodology and components. Section V entails an examination of the effectiveness of our method through a series of experiments conducted on various real-world networks. Section VI discusses the possible extension of MDEO. Lastly, Section VII concludes with a summary of our main results.
II Related work
@@
The capability of evolutionary optimization in handling discrete problems with non-linear characteristics has led to their extensive utilization within complex systems. For instance, Wang et al. [23] developed a multi-objective model by considering the nonuniform latency and computational complexity to enhance the tolerance of networks against attacks. Similarly, Wu et al. [24] leveraged the community distributions of networks to downsize the search space and introduce a multi-objective framework for network reconstruction by optimizing the reconstruction error and sparsity. Based on the same optimization objectives, Ying et al. [25] proposed a logistic principal component analysis-based method to improve the representation of networks. Moreover, evolutionary optimization has also been used in the diffusion source localization for sensor deployment to enhance the identification accuracy [26].
@@
Meanwhile, evolutionary algorithms were utilized to address the problem of community deception as well. Chen et al. [27] used the genetic algorithm to minimize the modularity and suggest the modifications to edges. Further, they categorized this problem into node communities, target communities, and all communities according to the scales [28]. Subsequently, a self-adaptive evolutionary framework is designed to streamline the search for the optimal edge set, complemented by the development of a permanence-based method to minimize the search space [29]. To mitigate the costs associated with evaluation, Zhao and Cheong [30] developed a divide-and-conquer strategy that partitions the network into sub-components, each optimized individually and cooperatively using the evolutionary algorithm. More applications of evolutionary algorithms on complex systems include dynamic community detection [31], and module identification [32]. These applications demonstrate the versatility and effectiveness of evolutionary algorithms in tackling diverse challenges within the field of complex network analysis.
@@
Recently, considerable efforts have been directed towards tackling evolutionary multi-task cases where the development of multi-factorial evolutionary algorithm [3] has laid the foundation for related studies. In addition, the optimization on single objective has been extended to the case of multiple objectives [33], adopting the implicit transfer mode where the crossover is operated on the solutions aiming at different tasks [34, 35]. In another extension of MTEO to the multi-objective setting [36], each task is decomposed into multiple subproblems, and the transfer probability of a solution is assessed based on how much it improves the performance of its corresponding subproblem. There is another route referred to explicit transfer that uses task-specific information to guide knowledge exchange between tasks by using statistical information [37] or constructing a mapping matrix [38, 39]. The primary focus of MTEO research lies in enhancing transfer strategies. For example, Wu et al. [40] proposed a diversified reasoning approach that expands solution space via varied transfer patterns. Xue et al. [41] used neural networks to evaluate task similarity and guide information flow, thereby enhancing transfer quality. In [42], reinforcement learning is used to learn effective policies where the knowledge transfer is modeled as sequential decisions.
TABLE I: Comparison of evolutionary optimization frameworks based on knowledge transfer. Multi-task evolutionary optimization is divided into two categories: MTEO-ConO, which addresses continuous problems, and MTEO-ComO, which focuses on combinatorial problems.
  Framework Domain Scope Problem Type Space Type Task Scope
MTEO-ConO [3, 33, 34, 35, 43, 44] Single-domain Continuous optimization Continuous Multiple
MTEO-ComO [21, 38, 45, 46, 47, 48, 49] Single-domain Combinatorial optimization Discrete Multiple
MDEO Multi-domain Combinatorial optimization Discrete Single/Multiple
@@
Beyond conventional areas, MTEO has been leveraged to strengthen the robustness of complex networks. For example, Wang et al. [22] utilized the correlation of tasks to develop a new crossover operator, thereby injecting knowledge into different individuals. According to [47], a graph neural network-based method has been developed that facilitates the exchange of information, demonstrating that structural destruction and cascading failure, though seemingly unrelated goals, can mutually enhance optimization outcomes. In a study on influence maximization [48], each task is formulated as a transformation, allowing for the effective application of MTEO and resulting in enhanced performance. Wu et al. [21] promoted the reconstruction of multiplex networks using MTEO with the inherent correlation of different layers. Similar research by Lyu et al. [50] formulates the modularity optimization within each layer as an independent task, facilitating community detection in multiplex networks.
Even though networks typically display common patterns, there has been minimal focus on cross-domain evolutionary optimization, especially with respect to network structures. Motivated by this overlooked potential, we introduce a novel optimization framework—multi-domain evolutionary optimization (MDEO). The comparison of the proposed MDEO and the existing MTEO literature is shown in Table I.
@@ III Optimization problem formulation
As a security-related task [51, 52, 53], community deception involves nodes intentionally concealing their true community affiliations by modifying (adding and removing) connections, which serves as the illustration of our MDEO for the following concerns:
Exploiting Community Detection Algorithms: With the assistance of community detection algorithms, densely interconnected groups of nodes with high intra-group connectivity and low inter-group connectivity can be easily identified [50, 54]. The development of community detection, while useful for understanding network structures and interactions, can also have certain dangerous aspects caused by the potential misuse or exploitation of community information [55, 56].
@@
Community deception serves as an innovative advancement in network science, radically altering node affiliations through minimal topological adjustments. Specifically, the modification of the network involves the addition of nonexistent edges and the deletion of existing edges. Suppose $G=\{V,E\}$ is a network where $V$ and $E$ denote the nodes and edges, respectively. The modified edges can be defined as follows:
[…]
where $E^{+}$ and $E^{-}$ denote the edges to add and delete, i.e.,
@@
where $\beta$ is the perturbance budget, referring to the maximum number of edges to rewire.
Assume there are two community structures belonging to the original network $G$ and modified network $G^{\prime}$, $\widetilde{\mathcal{C}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{k}\}$ and $\widetilde{\mathcal{C}^{\prime}}=\{\mathcal{C}^{\prime}_{1},\mathcal{C}^{\prime}_{2},\cdots,\mathcal{C}^{\prime}_{k^{\prime}}\}$, respectively, the problem of community deception is defined as:
@@
The function $\phi$ refers to the disparity between the community structures before and after modification. To validate the generalizability of MDEO, the node-level task is also tested in Section V-F.
IV Multi-Domain evolutionary optimization
This section provides a detailed introduction to the proposed MDEO, with its diagram shown in Figure 1. The mathematical notations used in MDEO are listed in Table II.
TABLE II: Mathematical notations.
  Symbols Definition
$G$ A network consisting of nodes $V$ and edges $E$
$\mathcal{G}$ Set of networks for optimization
$\widetilde{\mathcal{C}}$ Community distribution of the network $G$
$\beta$ Modification budget for the network $G$
$\mathcal{A}$ Set of anchor nodes in the networks
$v^{X,l}_{i}$ The $i$-th largest-degree node in the node set $X$
$v^{X,s}_{i}$ The $i$-th smallest-degree node in the node set $X$
$\Phi^{i\rightarrow j}$ Embedding mapping from network $G_{i}$ to network $G_{j}$
$M_{i\rightarrow j}$ Edge mapping from network $G_{i}$ to network $G_{j}$
$\mathcal{A}^{\widetilde{\mathcal{C}}}_{A,B}$ Aligned community between networks $G_{i}$ and $G_{j}$
$\mathcal{S}^{\mathcal{G}}_{i,j}$ Similarity bwtween networks $G_{i}$ and $G_{j}$
$T_{i\leftarrow j}$ Solution transferred from network $G_{j}$ to network $G_{i}$
Figure 1: The diagram of multi-domain evolutionary optimization. Four networks representing different systems (brain network, power network, social network, communication network) are optimized simultaneously. The transfer of knowledge across networks is achieved through the trained network alignment model. $e^{i}_{j}$ denote the $j$-th edge in $i$-th network and $M_{i\rightarrow j}$ is the edge mapping from the $i$-th network to the $j$-th network. The solution shown in yellow represents the transferred solution, while the light red color represents the elite solution from other networks. The red edge is the mutated edge that will be replaced by the edge of elite solutions of other networks. The knowledge exchange occurs when a series of conditions are met, and only elite solutions will be utilized to assist the optimization of other networks.